A batsman has a certain avrage of runs for 11 innings. In the 12 innings he scores 129 runs, there by increasing his average by 9 runs. What is his average for frist 11 innings?
Answer by Boreal(15235)

average unknown after 11 innings. Call it x
11x=total after 11 innings
11x+129 is total after 12 innings.
average is (11x+129)/12
That is 9 more than x
Therefore, (11x+129)/12=x+9
multiply everything by 12
11x+129=12x+108
x=21 runs per inning. ANSWER
for 11 innings 231 runs.
360 for 12 innings, average 30, increase of 9.